2. Average Senior High School annual cost of tuition fee for all private schools last year was Php43,700. A random sample of costs this year for 45 private schools indicated that the sample mean was Php45,800 and a sample standard deviation was Php5,600. At 0.01 level of significance, is there sufficient evidence to conclude that the cost increased? Solution: Step 1: State the hypothesis. 𝐻0: _____________________________________________________________ 𝐻1: _____________________________________________________________ Step 2: The level of significance and critical region. 𝛼 = ________ and the 𝑧𝑐𝑟𝑖𝑡𝑖𝑐𝑎𝑙 = __________ Step 3: Compute for the value of one sample t test. 𝑧𝑐𝑜𝑚𝑝𝑢𝑡𝑒𝑑 = _______ Step 4: Decision Rule. ________________________________________________________________ ___________________________________. Step 5: Conclusion: _______________


1
Expert's answer
2022-06-13T17:52:59-0400

Step 1: The following null and alternative hypotheses need to be tested:

"H_0:\mu\le43700"

"H_1:\mu>43700"

Step 2: This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=44" and the critical value for a right-tailed test is "t_c =2.414134."

The rejection region for this right-tailed test is "R = \{t:t>2.414134\}."

Step 3: The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{45800-43700}{5600/\sqrt{45}}=2.5156"


Step 4: Since it is observed that "t=2.5156>2.414134=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=44" degrees of freedom, "t=2.5156" is "p=0.007804," and since "p=0.007804<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Step 5: Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 43700, at the "\alpha = 0.01" significance level.